OpenStudy (anonymous):

You take a three-question true or false quiz. You guess on all the questions. What is the probability that you will get a perfect score?

OpenStudy (kirbykirby):

The probability of getting a correct answer on a given question is 1/2 To get all 3 questions correctly, then you must guess the 1st question correctly, as well as the 2nd and 3rd questions P(all 3 correct) = P(1st question correct)*P(2nd question correct)*P(3rd question correct) = (1/2)^3 = 1/8